OpenStudy (anonymous):

Check answer! I got for the answer y=3x+7. I need to know if that is correct. if not what the answer is please! The length of a rectangle on a coordinate plane has an equation y=2/5x + 7. Write the equation of a line on the coordinate plant that represents the equation of the width passing through the point (-6,2). y-y1=m(x-x1) Write answer in slope-intercept form:

OpenStudy (anonymous):

Please help if you can! for some reason i think i am doing it wrong, even though it seems simple. it is highschool geometry

OpenStudy (anonymous):

please if any of you can tell me the steps!

OpenStudy (anonymous):

They tell you the length is Y = 2/5 x +7..questions ask what is the line of the width on the point (-6,2). I would take the negative recripical slope of the length (because Width and Length are perpendicular lines) and use the point (-6,2) to determine the new line.

OpenStudy (anonymous):

neg recripical slope is -5/2 y - 2 = -5/2 (x+6) distribute the right side and as the question states...write in point slope formula

OpenStudy (anonymous):

oh, perpendicular. No wonder my picture looked like a parallelogram

OpenStudy (anonymous):

meant slope intercept form

OpenStudy (anonymous):

thank you so much :) ive been stuck on it over an hour

OpenStudy (anonymous):

"length of a rectangle on a coordinate plane"...per your question

OpenStudy (anonymous):

You are welcome

OpenStudy (anonymous):

yes, i realized that now thank you